OpenStudy (anonymous):

f(x)=3x+9 ,g(x)= 5x^2 find f(f+g)(x)

OpenStudy (anonymous):

f(f+g)(x) or f((f+g)(x))???

OpenStudy (anonymous):

F(f+g)(x)

OpenStudy (anonymous):

first do (f+g)(x) that is (f+g)(x) = f(x) +g(x) now you will get some eqaution in x after doing the above step lets call it P then P = (f+g)(x) = f(x) +g(x) now all you have to do is f(P)